OpenStudy (anonymous):

change these equations into slope intercept form -2x-1=y x=-y+3 x+y=3 PLEASE HELP WITH ALL!!!!!!!

OpenStudy (anonymous):

1. its still y= but flip the -2x and th -1 so its y-=-2x-1 2. solve for y 3. solve for y

OpenStudy (anonymous):

y=-2x-1 is already in slope intercept form. slope intercept form is y=mx + c and @blondie16 is right for the last two.

OpenStudy (anonymous):

so would the second one be y=3-x and the last one be 3-x=y @blondie16

OpenStudy (anonymous):

put them in the form of y=mx+b like @JayDS said

OpenStudy (anonymous):

y=x-3 and the same for the last one

OpenStudy (anonymous):

@blondie16

OpenStudy (anonymous):

the first one would be y=-x+3 the other one would be y=-x+3 too

OpenStudy (anonymous):

ok thankss

OpenStudy (anonymous):

medal?